In a probability experiment, Karen flipped a coin 54 times. The coin landed on heads 33 times. What percentage of the coin flips resulted in tails? Round to the nearest percent.
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

In a probability experiment, Karen flipped a coin 54 times. The coin landed on heads 33 times. What percentage of the coin flips resulted in tails? Round to the nearest percent.
-----------------
Solve for x:
x/100 = 33/54
---
x = 100(33/54)
x = 61.11%
===============
Cheers,
Stan H.

Answer by ewatrrr(24785)


Hi
Karen flipped a coin 54 times. The coin landed on heads 33 times.
What percentage of the coin flips resulted in tails?
(54-33)/54 = .389 OR 39% (rounded to the nearest percent)